Demo file for DVIWIN 2.8 September 30, 1993 Hippocrates Sendoukas

This driver lets you preview and print DVI files under MS-Windows 3.1. Its main advantages are: speed, compatibility with any raster device with a Windows driver, and graphics capability. Apart from some assembly acceleration for all machines, it uses special routines on 80386 or 80486 processors to speed up the rasterization. All screen and printer handling is done through Windows, so it should work on any printer supported by the system. You can insert arbitrary graphics files produced by most Windows applications, or other standard graphics files. This version has intrinsic support for WMF, BMP, PCX and MSP files; other formats (eg., TIFF, HPGL, etc.) can also be imported provided that you have the appropriate graphics filter. The following graph was made with Excel and it will come out in color if you print it on a color printer.

Sample Figure: Utility surface for two goods

The driver also understands emTeX and tpic specials. The box and the shading of this paragraph was made with a few simple tpic specials.

For more information on using the programs, please read or print the files “dviwin.wri”, “wbr.wri” and “clipmeta.wri”. If you encounter any installation or operation problems, read the file “helpme.wri” and if it does not resolve the problem, please let me know. My address is “3 Sifnou St., Athens 11254, Greece”